OpenStudy (ilovecake):

Evaluate the expression when x = 6. 3x + 7 A. 25 B. 39 C. 43 D. 126

OpenStudy (eric_d):

Substitute x=6

OpenStudy (eric_d):

into that equation @Ilovecake

OpenStudy (ilovecake):

Ok thanks :)

OpenStudy (eric_d):

You're most welcome @Ilovecake

OpenStudy (ilovecake):

@eric_d that is all I need to do?

OpenStudy (eric_d):

Yes

OpenStudy (eric_d):

So, what's the answer

OpenStudy (ilovecake):

@eric_d Ok thanks again

OpenStudy (eric_d):

Try nw

OpenStudy (eric_d):

@Ilovecake

OpenStudy (eric_d):



OpenStudy (ilovecake):

6 - 2 = 4 and 1- (-2)+(-2)2 = 1+2+4=7........right?

OpenStudy (eric_d):